# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 3.3.2 Amperometry
Method Form Factor Targets Analytes Evaluation Control Results Notes Source
Amperometry Benchtop Vitamin B9 (folic acid) Serum Human subjects Spiked known concentrations >99% recovery of spiked concentration Robust against interference from ascorbic acid; sensor performance peaked at pH 7.2, slightly lower than pH of serum; folic acid is the synthetic form of folate [106]
Amperometry Wearable tattoo Vitamin C (ascorbic acid) Sweat, tears, and saliva Human subjects Supplementation Similar time dynamics to plasma for tears and sweat [151]
Amperometry Wearable sensor chip Vitamin C (ascorbic acid) Sweat, urine, and blood Human subjects Supplementation/ intake, blood measurements with the same sensor Urine and sweat measurements increased after intake, with inter-trial variation. Correlations with blood sensor measurements were 0.81 for sweat and 0.72 for urine Used emergen-C as a vitamin C supplement, which contains several other micronutrients [190]
Electrocatalytic vitamin C amperometric assay and competitive vitamin D immunoassay PoC Vitamins C (ascorbic acid) and D (calcidiol) Saliva Human subjects Vitamin supplementation Observable rise and drop in levels over time No quantitative results; saliva sample must be pre-treated; Vitamin C has little evaluation [145]
Amperometry Wearable sensor chip Vitamin C (ascorbic acid) Sweat and saliva Stock solutions and human subjects Known solutions and supplementation $R^{2}$ of 0.99 and LoD of 0.0299 mcg/mL for stock solutions, was able to detect general increase in current after supplementation [191]
Chronoamperometry Wearable sensing system Vitamin C (ascorbic acid) and Levodopa Sweat Human subjects Vitamin supplementation Detection correlation coefficients of both exceed 0.99; both sensors have a wide linear detection range of 0-17.6 mg/L and 0-1000 $\mu$M, respectively, and low detection limits of 0.05 mg/L and 17.9 $\mu$M, respectively. The system is wireless, battery-free, flexible, finger-actuated, and self-pumping [122]
3.3.3 Others
Two techniques do not fit neatly into the above categories: body impedance measurement scanning [65] and electroacupuncture [3]. Heo et al. [65] focuses on analyzing solely vitamin D (calcidiol) status using body impedance. They explored the correlation between vitamin D levels in blood, body composition, blood parameters from checkup, and arm impedance (from wrist to elbow) to calibrate an impedance measurement frequency for vitamin D [65]. The motivation was that "body fat accumulates vitamin D," and body fat can be measured by impedance measurement [65].
Vitastiq is the most immediately relevant device for micronutrition insights, with the company claiming that the device can non-invasively measure 26 different vitamins and minerals [3]. The basis of the device’s technology is electroacupuncture. It measures the electrical resistance at an ’acupuncture point’ and compares it to a ’calibration point’. These points are very specific, and the user is taken through a tutorial to locate them properly [92]. Minimal details about how the device functions and electroacupuncture provide insight into micronutritional status are available. The information the user receives is only qualitative, and consumer reviewers noted a lack of confidence in the results [92]. As such, the manufacturer notes that Vitastiq is not a medical device and should only be used as an indicator of "general vitamin trend" [3].
Table 6: Other Electrochemistry-Based Methods
@@ 3.3.3 Others
Method Form Factor Targets Analytes Evaluation Control Results Notes Source
Arm impedance measurement scan Non-mobile, clinical device Vitamin D (calcidiol) Impedance measurement at 21.1 Hz Human subjects Vitamin D status by blood test (method unspecified) $R^{2}$ of 0.75 (regression with Vitamin D level) Also evaluated regression models with medical checkup and body composition analysis data [65]
Electroacupuncture PoC 26 different vitamins and minerals Electrical resistance on the surface of the skin Not published Not published Not published Should only be used as an indicator of "general vitamin trend" [3]
3.4 Spectroscopy-Based Methods
Spectroscopy is the most common method of gold-standard micronutrient status assessment. However, there is still a lot of work to do to make spectroscopic approaches more accessible and less invasive. We begin with some definitions of common terms in the field. Spectroscopy is "the investigation and measurement of spectra produced by matter interacting with or emitting electromagnetic radiation" [2]. Spectrometry is the application of spectroscopy; the way in which quantitative measurements are obtained. When we speak of a spectra, it means any measurement that is a function of wavelength or frequency. A sample will absorb or emit these spectra when electromagnetic radiation of a known wavelength is applied. Spectra are measured by a detector, the spectrometer. Because the level of radiation applied is known, analyzing the resultant spectra after it interacts with the sample provides information about the sample.
We have previously described MS and LC-coupled spectroscopy, the types of spectroscopy used by most gold-standard clinical biochemical analyses (Section 3.1. As mentioned, these methods are expensive, non-specific, complex, and often require an invasively-collected biosample. While there have been strides to make MS more accessible [35], we focus on alternative spectroscopic techniques that may yield micronutrient insights.
Categorized under emission spectroscopy, fluorescence spectroscopy has been demonstrated for the measurement of primarily B vitamins in non-biological samples such as multivitamins and energy drinks, although vitamin B1 was measured in urine [189]. The B vitamins continue to get attention in near infrared (NIR) spectrophotometry (750-2500 nm wavelength), where their measurement has been reported as well [189]. One review recognizes the ability of vibrational spectroscopy, which includes infrared (IR) and raman spectroscopy, to act as a tool for biofluid analysis in precision nutrition [40] (Fig. 3). However, like other studies in nutrition, this review has few considerations for micronutrients and the approaches covered are largely concerned with general nutritional status or macronutrients. Tsiminis et al. [170] notes the potential of raman spectroscopy for measuring vitamin B12, though this has yet to be realized at in-body concentrations due to the low sensitivity of raman spectroscopy. In biosamples, spectrophotometry was applied to measure vitamin C [69]. Spectroscopic skin tests and raman spectroscopy have also been noted as promising techniques in the assessment of provitamin A carotenoid status in the body [69]. Measurement of several water-soluble vitamins in synthetic mixtures and dosage forms was achieved with derivative and multivariate spectrophotometry [189]. Derivatives of UV spectrophotometry (185-400 nm wavelength) have also been particularly useful in the analysis of caffeine and B vitamins in energy drinks.
Figure 3: Vibrational spectroscopy targeted at saliva, tissue, skin, urine, and stools holds potential nutritional insights. Used with permission from [40]
3.4.1 Accessible Spectroscopy
Significant progress has been made to make spectrometry as a study more accessible and compact [35]. Major subfields of spectrometry (visible, Raman, mid-IR, NIR, MS, and hyperspectral imaging) have seen the development of portable or handheld devices. In some cases, such as visible, near-IR, and hyperspectral imaging, these can even be smartphone-based. This paves the way for more accessible, noninvasive techniques. A summary of spectroscopy-based methods is found in Table 7.
There was one application of spectrometry for vitamin D (calcidiol) measurement in our interest area of accessible approaches [178]. With human serum samples in mind, their sensor used surface plasmon resonance (SPR) together with smartphone-based spectrophotometry to assess vitamin D content. The general design of the sensor involved an optical waveguide to direct light from the smartphone flash through one or more SPR sensors, a diffraction grating, and finally into the smartphone camera where a spectra of pixel intensities can be extracted. The device was evaluated on spiked serum samples. As the concentration decreased, the center of mass of the spectra shifted right, allowing for the detection of these concentrations. The paper claims a comparable LoD to gold standard methods of LC-MS, but no quantitative estimates of vitamin D concentration or statistical analysis of the results are provided.
Some benchtop approaches to spectroscopic analysis of biofluids for micronutrient assessment have also been demonstrated [132, 133, 19, 116]. Peterson et al. [132] present a photonic crystal-based sensor for ferritin assessment. Photonic crystals are designed to accumulate a target biomolecule on their surface, which changes their reflected peak wavelength value (PWV) under a spectrophotometer. The authors subjected their sensor to robust evaluation, utilizing human liver ferritin, commercial serum controls (Liquichek), and three different ELISA tests. The developed sensor held up against the ELISA tests, but with a higher LOD that doesn’t cover the lower end of deficiency, into anemia (26 mcg/L). Bias by Bland-Altman analysis was similar to that of the BioVendor ELISA, and recovery from known serum controls was >94%.
The same first author identifies a specificity issue with their previous assay [132] and employs iron-oxide nanoparticles to minimize non-specific signals [133]. This time, the goal was soluble transferrin receptor (sTfR) quantification from serum, an indicator of iron supply to tissues [6]. Biomolecule interaction on the assay was quantified using the Biomolecular Interaction Detection system from SRU Biosystems Inc. The authors compared results from their assay to ELISA and the previously developed photonic crystal assay. The authors claim that the bias of the assay was not “statistically different from the reference ELISA tests” [133].
Moving from iron to vitamin C, Bi et al. [19] demonstrates the immobilization of ascorbate oxidase in a microfluidic channel, enabling the quantification of vitamin C with UV-visible spectroscopy. During analysis, the biosample is diluted in phosphate buffered saline (PBS) and “pumped through the microfluidic channel”. A serum sample was obtained by a single healthy, female volunteer for evaluation. The sample was pretreated to remove proteins , and a few drops were added to the sensor. Even with extensive pretreatment, there was evidence of interference at 280 nm, close to the analysis peak of vitamin C at 266 nm. No gold standard measurement was provided for comparison. Mughal et al. [116] mix different electrolytic solutions with plasma and serum, and when paired with a novel reduced graphene oxide, vitamins K1, K2, B6, and D3 can be individually identified using UV-visible spectrophotometry. Though this method did not compare measured levels with clinical values.
Table 7: Spectroscopic Methods
@@ 3.4.1 Accessible Spectroscopy
Method Form Factor Targets Analytes Evaluation Control Results Notes Source

### v2

@@ 3.3.1. Voltammetry
Method Platform Targets Analytes Evaluation Control Results Notes Source
Voltammetry Benchtop Vitamins B2 (riboflavin), B9 (folic acid), and C (ascorbic acid) Plasma Human subjects Spiked known concentrations ¿99% recovery of spiked concentrations Tested in 7.2 pH buffer, which is slightly lower than pH of blood ($\sim$7.4); direct riboflavin is not the gold-standard biomarker for vitamin B2; folic acid is the synthetic form of folate; Linear range for vitamin C did not reach below the lower limit for sufficiency (Revin and John, 2012)
Anodic stripping voltammetry Benchtop Zinc Serum Human subjects Spiked known concentrations Peak current decreased with concentration, but were lower in magnitude than buffer A sample pH of 6 was necessary for optimal performance (Jothimuthu et al., 2013)
SWV Benchtop Vitamin B9 (folic acid), GSH, and NADH Urine and serum Human subjects (urine only) Spiked known concentrations Accurate recovery of spiked concentration Simultaneous determination of targets; unclear whether pre-existing urine composition biased recovery; claims serum evaluation but this is not presented; folic acid is the synthetic form of folate (Raoof et al., 2015)
Square wave anodic stripping voltammetry Wearable tattoo Zinc Sweat Zinc stock solutions Known zinc solutions $R^{2}$ of 0.999 for measured current vs stock solutions, LoD of 0.05 $\mu$g/mL Zinc content in actual sweat from single participant was close to physiological range (Kim et al., 2015a)
Square wave anodic stripping voltammetry Wearable patch Zn, Cd, Pb, Cu, and Hg ions Sweat and urine Human subjects ICP-MS Similar to control, but provided no statistical analysis Included temperature sensor to account for the influence of skin temperature on peak current (Gao et al., 2016)
SWV Benchtop Vitamin B12 (cyanocobalamin) Urine Diluted from human subjects Spiked known concentrations 98-105% recovery Cyanocobalamin is the synthetic form of B12 (Stanković et al., 2016)
SWV and chronoamperometry Eyeglasses Vitamins B2, B6, and C, alcohol, and glucose Tears Human subjects Breathalyzer BAC, commercial glucometer, vitamin supplementation Correlations of 0.852 with BAC,  0.7 with glucometer, distinct voltage peaks for each vitamin Glucose and alcohol was main focus, vitamin assessment included as a proof of concept; exact form of each vitamin is not known (Sempionatto et al., 2019)
Cyclic voltammetry Benchtop Iron (ferritin) Serum Spiked from human subjects ELISA $R^{2}$ of 0.966 and lower linearity than control; tended to overestimate concentrations ¡100 $\mu$g/L Range of spiked concentrations was not sensitive to deficiency; studied the impact of pH and interferent compounds (Garg et al., 2020)
Cyclic voltammetry Portable Vitamin C Serum Human subjects HPLC $R^{2}$ of 0.984 (p¡0.001) and 100% accuracy in deficiency detection Targets scurvy (extreme deficiency) but the threshold was set above sufficiency (Sun et al., 2021)
Voltammetry Wearable patch Vitamins B6 (pyridoxine), C (ascorbic acid), D3 (calciol), E (alpha-Tocopherol), and other macronutrients and amino acids Sweat Not reported for vitamins Not reported for vitamins Observable linear relationship between vitamin concentration and peak height current density Only vitamin C aligns with clinical standard, quantitative results for vitamins were not reported (Wang et al., 2022)
Voltammetry Benchtop Vitamin D3 Plasma Spiked from human subjects Spiked known concentrations and HPLC-UV LoD of 1.9 ng/mL; RSD of 0.3-2.6% and recovery of 96-102% Exact form of D3 (gold-standard 25(OH)D3/calcidiol or calciol) not reported (Lokesh Kumar et al., 2023)
SWV and potentiometric measurement Portable Zinc and vitamin C (ascorbic acid) Fingertip sweat Human subjects Supplementation Both micros could be analyzed over time simultaneously No comparison to clinical assessment of status or statistical analysis of results; Vitamin C range far exceeded physiological concentrations in plasma (Seker et al., 2024)
Differential pulse voltammetry Wearable patch Vitamin B2 (riboflavin) Sweat Human subjects HPLC $R^{2}$ of 0.9783 with sweat HPLC and 0.87 with urine fluorescence spectroscopy Urine included as comparison to sweat status; incorporates pH sensor to control for influence of pH on measurements; direct riboflavin is not the gold-standard biomarker for Vitamin B2 status (Shi et al., 2024)
3.3.2. Amperometry
Vitamins B2, B9, C, and D, as well as the mineral iron, have been assessed in biofluids using amperometry (Maiyalagan et al., 2013; Sempionatto et al., 2020; Zhao et al., 2021; Ruiz-Valdepeñas Montiel et al., 2021; Zhu et al., 2023; Ning et al., 2024). Maiyalagan et al. (2013) confronted a major limitation of glassy carbon electrode-based sensors for vitamin B9 (folic acid): interference from vitamin C. Their nanofiber-modified electrode successfully avoided this interference. In evaluation, two serum samples were collected and evaluated before and after spiking with 4.41 $\mu$g/L of folic acid. The peak current increased accordingly, allowing for greater than 99% recovery. Vitamin C is examined by Sempionatto et al. (2020), who deployed amperometry and immobilized ascorbate oxidase in a tattoo-based platform. Tests on human subjects (n=4) focused on the temporal characteristics of the current response, finding peak response in sweat 90 minutes after supplementation and a return to baseline 180 minutes after, in line with the plasma response of vitamin C. Tears and saliva were noted as other possible biofluids, with tears yielding a similar temporal profile on a single subject (albeit with different peak currents). The authors also experimented with supplementing vitamin C through orange juice, claiming that the response from sweat samples increased in line with increasing vitamin C content (n=2). Crucially, no statistical analysis was conducted on the results of the experiments, such as correlations between intake and measured current. There were also no comparisons made to clinical reference methods of vitamin C assessment, though this was stated as a subject for future work.
A wearable, electrochemical device to measure vitamin C levels in sweat, urine, and blood was proposed by Zhao et al. (2021). The device was used in a study where 6 male participants, aged 20-30, were given vitamin C as emergen-C brand supplements. The sensor was wearable, but no on-body measurements were made. Instead, urine and induced sweat were collected three hours post-intake and measured with the device. Blood samples were collected from a single participant in a separate study, analyzed with the device, and compared to results from urine and sweat. This research did not compare device results to any optimal reference standard for vitamin C assessment; instead, it relied on intake, which (as we will see in Section 4) is a poor equivalent for in-body status due to individual differences in micronutrient absorption. In addition, we note that the vitamin C supplement, emergen-C, contains several other nutrients that could influence the results of the analysis.
One paper proposed simultaneous measurement of vitamins C and D from a single saliva sample (Ruiz-Valdepeñas Montiel et al., 2021). Their sensor combined an electrocatalytic vitamin C (ascorbic acid) amperometric assay and competitive vitamin D (25(OH)D3) immunoassay. Vitamin D was the clear focus of the paper, and the vitamin C sensor received little to no attention aside from some analysis of potential cross-talk between the two sensors. The sensor was applied in a study that supplemented vitamins to 3 participants and used the device to analyze saliva samples at increasing time intervals from intake. No optimal reference standard assessment method was used to evaluate the sensors, although the authors advocated for this evaluation and the development of truly quantitative sensors in future work. Another flexible, electrochemical sweat biosensor for vitamin C used polyaniline film modified with phytic acid (Zhu et al., 2023). The biosensor was validated with synthesized vitamin C samples of known concentrations. Four human subjects were given supplements and had their sweat collected 3 times over 90 minutes. The sensor detected a general increase in current from the sweat samples over time, with variation across subjects. Saliva tests were also conducted. Peak current occurred 60 minutes after supplementation, in line with results from (Sempionatto et al., 2020). Another team of researchers designed a finger-actuated wirelessly-charging wearable that measures vitamin C and levadopa (a central nervous system agent) levels from sweat (Ning et al., 2024). The system had a microfluidic chip with a self-driven pump and anti-reflux valve, a flexible wireless circuit board, and a companion smartphone app. They ran a study with five healthy participants whose sweat was collected and measured after exercising and ingesting vitamin C tablets as well as fava beans (Ning et al., 2024). The results were not compared to clinical results as a optimal reference standard baseline.
Table 6. Electrochemistry-Based Methods Using Amperometry
@@ 3.3.2. Amperometry
Method Platform Targets Analytes Evaluation Control Results Notes Source
Amperometry Benchtop Vitamin B9 (folic acid) Serum Human subjects Spiked known concentrations ¿99% recovery of spiked concentration Robust against interference from ascorbic acid; sensor performance peaked at pH 7.2, slightly lower than pH of serum; folic acid is the synthetic form of folate (Maiyalagan et al., 2013)
Amperometry Wearable tattoo Vitamin C (ascorbic acid) Sweat, tears, and saliva Human subjects Supplementation Similar time dynamics to plasma for tears and sweat (Sempionatto et al., 2020)
Amperometry Wearable sensor chip Vitamin C (ascorbic acid) Sweat, urine, and blood Human subjects Supplementation/ intake, blood measurements with the same sensor Urine and sweat measurements increased after intake, with inter-trial variation. Correlations with blood sensor measurements were 0.81 for sweat and 0.72 for urine Used emergen-C as a vitamin C supplement, which contains several other micronutrients (Zhao et al., 2021)
Amperometry Wearable sensor chip Vitamin C (ascorbic acid) Sweat and saliva Stock solutions and human subjects Known solutions and supplementation $R^{2}$ of 0.99 and LoD of 0.0299 $\mu$g/mL for stock solutions, was able to detect general increase in current after supplementation (Zhu et al., 2023)
Electrocatalytic vitamin C amperometric assay and competitive vitamin D immunoassay Portable Vitamins C (ascorbic acid) and D (calcidiol) Saliva Human subjects Vitamin supplementation Observable rise and drop in levels over time No quantitative results; saliva sample must be pre-treated; Vitamin C has little evaluation (Ruiz-Valdepeñas Montiel et al., 2021)
Chronoamperometry Wearable sensing system Vitamin C (ascorbic acid) and Levodopa Sweat Human subjects Vitamin supplementation Detection correlation coefficients of both exceed 0.99; both sensors have a wide linear detection range of 0-17.6 mg/L and 0-1000 $\mu$M, respectively, and low detection limits of 0.05 mg/L and 17.9 $\mu$M, respectively. The system is wireless, battery-free, flexible, finger-actuated, and self-pumping (Ning et al., 2024)
3.3.3. Impedance Analysis
Heo et al. (2021) focused on analyzing solely vitamin D (calcidiol) status using body impedance. They explored the correlation between vitamin D levels in blood, body composition, blood parameters from checkup, and arm impedance (from wrist to elbow) in 26 patients (14 male) to calibrate an impedance measurement frequency for vitamin D (Heo et al., 2021). The motivation was that ”body fat accumulates vitamin D,” and body fat can be measured by impedance measurement (Heo et al., 2021).
Table 7. Other Electrochemistry-Based Methods
@@ 3.3.3. Impedance Analysis
Method Platform Targets Analytes Evaluation Control Results Notes Source
Arm impedance measurement scan Non-mobile, clinical device Vitamin D (calcidiol) Impedance measurement at 21.1 Hz Human subjects Vitamin D status by blood test (method unspecified) $R^{2}$ of 0.75 (regression with Vitamin D level) Also evaluated regression models with medical checkup and body composition analysis data (Heo et al., 2021)
3.3.4. Comparison
Across voltammetry and amperometry, vitamin C, B vitamins, and zinc are the most common targets of electrochemistry-based methods. However, a few do target vitamin D, which is the only target of the impedance analysis technique. Voltametric devices were more often applied to blood or urine biosamples, and the associated studies relied less on supplementation as a control compared to amperometry-based methods. Amperometric devices appeared to focus more on a wearable form-factor. The single study using impedance analysis reported results which were less accurate than those reported by voltametric or amperometric approaches to vitamin D assessment (Heo et al., 2021). This may indicate that body impedance analysis is not a promising method for future research, but we add that the results of this study may be more trustworthy than others due to its larger sample size and evaluation in the context of optimal reference standard vitamin D status assessments using blood. For these reasons, we cannot conclude whether impedance analysis is less promising than amperometry or voltammetry.
3.4. Spectroscopy-Based Methods
Spectroscopy is the most common method of optimal reference standard micronutrient status assessment. However, there is still a lot of work to do to make spectroscopic approaches more accessible and less invasive. We begin with some definitions of common terms in the field. Spectroscopy is ”the investigation and measurement of spectra produced by matter interacting with or emitting electromagnetic radiation” (noa, 2019). Spectrometry is the application of spectroscopy; the way in which quantitative measurements are obtained. When we speak of a spectra, we mean any measurement that is a function of wavelength or frequency. A sample will absorb or emit these spectra when electromagnetic radiation of a known wavelength is applied. Spectra are measured by a detector, the spectrometer. Because the level of radiation applied is known, analyzing the resultant spectra after it interacts with the sample provides information about the sample.
We have previously described MS and LC-coupled spectroscopy, the types of spectroscopy used by most optimal reference standard clinical biochemical analyses (Section 3.1. As mentioned, these methods are expensive, non-specific, complex, and often require an invasively-collected biosample. While there have been strides to make MS more accessible (Crocombe, 2018), we focus on alternative spectroscopic techniques that may yield micronutrient insights.
Categorized under emission spectroscopy, fluorescence spectroscopy has been demonstrated for the measurement of primarily B vitamins in non-biological samples such as multivitamins and energy drinks, although vitamin B1 was measured in urine (Zhang et al., 2018). The B vitamins continue to get attention in near infrared (NIR) spectrophotometry (750-2500 nm wavelength), where their measurement has been reported as well (Zhang et al., 2018). One review recognized the ability of vibrational spectroscopy, which includes infrared (IR) and raman spectroscopy, to act as a tool for biofluid analysis in precision nutrition (Dongdong and Cozzolino, 2023) (Fig. 3). However, like other studies in nutrition, this review had few considerations for micronutrients and the approaches covered were largely concerned with general nutritional status or macronutrients. Tsiminis et al. (2017) noted the potential of raman spectroscopy for the measurement vitamin B12, though this has yet to be realized at in-body concentrations due to the low sensitivity of raman spectroscopy. In biosamples, spectrophotometry was applied to measure vitamin C (Höller et al., 2018). Spectroscopic skin tests and raman spectroscopy have also been noted as promising techniques in the assessment of provitamin A carotenoid status in the body (Höller et al., 2018). Measurement of several water-soluble vitamins in synthetic mixtures and dosage forms was achieved with derivative and multivariate spectrophotometry (Zhang et al., 2018). Derivatives of UV spectrophotometry (185-400 nm wavelength) have also been particularly useful in the analysis of caffeine and B vitamins in energy drinks. The open-source Lumos platform also enables on-body spectroscopy (Watson et al., 2022).
Figure 3. Vibrational spectroscopy targeted at saliva, tissue, skin, and urine have current applications in precision nutrition, while stools hold potential nutritional insights (indicated by question marks). Used with permission from Dongdong and Cozzolino (2023)
[…]
The figure shows a clipart outline of the human body with five categories labeled: Saliva, Tissue, Skin, Urine, and Stools. Each of these categories has brackets detailing the specific nutritional insights that can be gained from each sample. Saliva provides information on food interactions and saliva biomechanical and chemical composition, tissue lends information about body fat, lean/fat, and BMI, skin provides information on antioxidants and carotenoids, urine shows nutrients and biomarkers, and stools provide insights on the microbiome, digestibility, and intake.
3.4.1. Accessible Spectroscopy
Significant progress has been made to make spectrometry as a study more accessible and compact (Crocombe, 2018). Major subfields of spectrometry (visible, Raman, mid-IR, NIR, MS, and hyperspectral imaging) have seen the development of portable or handheld devices. In some cases, such as visible, near-IR, and hyperspectral imaging, these can even be smartphone-based. This paves the way for more accessible, noninvasive techniques. A summary of spectroscopy-based methods is found in Table 8.
There was one application of spectrometry for vitamin D (calcidiol) measurement in our interest area of accessible approaches (Walter et al., 2020). With human serum samples in mind, their sensor used surface plasmon resonance (SPR) together with smartphone-based spectrophotometry to assess vitamin D content. The general design of the sensor involved an optical waveguide to direct light from the smartphone flash through one or more SPR sensors, a diffraction grating, and finally into the smartphone camera where a spectra of pixel intensities can be extracted. The device was evaluated on spiked serum samples. As the concentration decreased, the center of mass of the spectra shifted right, allowing for the detection of these concentrations. The paper claims a comparable LoD to optimal reference standard methods of LC-MS, but no quantitative estimates of vitamin D concentration, sample size, or statistical analysis of the results are provided.
Some benchtop approaches to spectroscopic analysis of biofluids for micronutrient assessment have also been demonstrated (Peterson et al., 2014, 2015; Bi et al., 2016; Mughal et al., 2024). Peterson et al. (2014) presented a photonic crystal-based sensor for ferritin assessment. Photonic crystals were designed to accumulate a target biomolecule on their surface, which changes their reflected peak wavelength value (PWV) under a spectrophotometer. The authors subjected their sensor to robust evaluation, utilizing human liver ferritin, commercial serum controls (Liquichek), and three different ELISA tests. The developed sensor held up against the ELISA tests, but with a higher LoD (26 $\mu$g/L) that did not cover the lower end of iron deficiency. Bias by Bland-Altman analysis was similar to that of the BioVendor ELISA, and recovery from known serum controls was greater than 94%.
Motivated by a specificity issue in the preceding sensor, Peterson et al. (2015) employed iron-oxide nanoparticles to minimize non-specific signals. This time, the goal was soluble transferrin receptor (sTfR) quantification from serum, an indicator of iron supply to tissues (Aggett, 2012). It should be noted that sTfR is not influenced by inflammation to the same extent as serum ferritin, and is therefore a potentially more ideal biomarker. However, sTfR is more expensive to assay. For this experiment, biomolecule interaction on the assay was quantified using the Biomolecular Interaction Detection system from SRU Biosystems Inc (Peterson et al., 2015). The authors compared results from their assay on Liquichek control sera to ELISA and the previously developed photonic crystal assay. The authors claimed that the bias of the assay was not “statistically different from the reference ELISA tests” (Peterson et al., 2015).
Moving from iron to vitamin C, Bi et al. (2016) demonstrated the immobilization of ascorbate oxidase in a microfluidic channel, enabling the quantification of vitamin C with UV-visible spectroscopy. During analysis, the biosample was diluted in phosphate buffered saline (PBS) and “pumped through the microfluidic channel”. A serum sample was obtained by a single healthy, female volunteer for evaluation. The sample was pretreated to remove proteins , and a few drops were added to the sensor. Even with extensive pretreatment, there was evidence of interference at 280 nm, close to the analysis peak of vitamin C at 266 nm. No optimal reference standard measurement was provided for comparison. Mughal et al. (2024) mixed different electrolytic solutions with plasma and serum, and when paired with a novel, reduced graphene oxide, vitamins K1, K2, B6, and D3 could be individually identified using UV-visible spectrophotometry. This method did not compare the measured levels from 5 subjects with clinical values.
Table 8. Spectroscopic Methods
Method Platform Targets Analytes Evaluation Control Results Notes Source